Thank you in advance!

Jill is 8 years older than her brother Curt. Four years from now, Jill will be twice as old as Curt. How old is each at the present time?

Answer by ankor@dixie-net.com(22740)

Jill is 8 years older than her brother Curt. Four years from now, Jill will be twice as old as Curt. How old is each at the present time?
:
J = Jill's present age; C = Curt's present age
:
The equation for this statement: "Jill is 8 years older than her brother Curt."
J = (C + 8)
:
"Four years from now, Jill will be twice as old as Curt."
(J + 4) = 2(C + 4)
J + 4 = 2C + 8
:
Substitute (C + 8) for J
(C+8) + 4 = 2C + 8
C + 12 = 2C + 8
C - 2C = 8 - 12
-C = -4
C = 4 yrs old now
:
That make Jill; 12 yrs old now
:
In 4 years: Jill will be 16, twice as old as Curt, who will be 8